# Preference Curriculum: LLMs Should Always Be Pretrained on Their Preferred Data

### v1

@@ Abstract
Current large language models (LLMs) generally utilize a consistent data distribution throughout the entire pretraining process. However, as the model’s ability improves, it intuitively should be pretrained with differentiated data. To achieve it, we propose the Perplexity Difference based Preference Curriculum learning (PDPC) framework, which always perceives and uses the data preferred by LLMs to train and boost them. Firstly, we introduce the PD metric to measure the difference in how well strong and weak models fit the samples. Samples with high PD are more challenging for weak models to learn and are more suitable to be arranged in the later stage of pretraining. Secondly, we propose the PD preference function to approximate the model and predict the data preference of the LLM at any time, so as to complete the arrangement of the entire data offline and ensure continuous training without interruption. Experimental results on 1.3B and 3B models demonstrate that our PDPC significantly surpasses baselines. Notably, the 3B model achieved more substantial gains, with an increased average accuracy of over 4.1% across various benchmarks.
@@
Xuemiao Zhang1,4,  Liangyu Xu4,  Feiyu Duan2,4thanks: Equal contribution.,  Yongwei Zhou4,  Sirui Wang3,4thanks: Corresponding author.,  Jingang Wang4,  Xunliang Cai4 1 Peking University 2 Beihang University 3 Tsinghua University 4 Meituan zhangxuemiao@pku.edu.cnduanfeiyu@buaa.edu.cnywzhouphd2018@gmail.com {xuliangyu02, wangsirui, wangjingang02, caixunliang}@meituan.com
@@ 1 Introduction
Large language models (LLMs) have shown impressive performance across a variety of tasks after being pretrained on vast amounts of data Touvron et al. (2023); Dubey et al. (2024); Liu et al. (2024). As LLMs undergo extensive pretraining, their capabilities steadily improve, which influences their performance, characteristics, and data preferences Yu et al. (2024). Existing methods of uniformly sampling data during the whole pretraining process are suboptimal because they overlook the model’s evolving data preferencesWettig et al. (2024); Abbas et al. (2023); Sachdeva et al. (2024).
Recently, some research has shifted the focus to the influence of data on model capability during the pretraining processEvans et al. (2024); Yu et al. (2024); Koh and Liang (2017); Ko et al. (2024). A typical method named MATES considers the changing data influence on models but requires interrupting the training process to select more preferred data based on the model’s current state, which disrupts training continuity and stability, ultimately reducing pretraining efficiency Yu et al. (2024). A similar issue arises in JEST Evans et al. (2024).
In this paper, we propose perplexity Ziegel et al. (1976) difference (PD) to measure the difference in the fit of LLMs’ final and early checkpoints. We further propose a novel PD-based Preference Curriculum learning framework, PDPC. It’s required to perceive LLMs’ data preference at any training step and use the preferred data to continuously pretrain LLMs without interruption, thereby boosting their performance. To this end, PDPC needs to solve three major challenges:
@@
Ideally, training would be paused at any point to calculate PD using the model’s current state, allowing for adjustment of the preference distribution like MATESYu et al. (2024). However, to ensure continuity and stability in pretraining, we propose an offline processing method to approximate the ideal dynamic preference adjustment. First, we calculate PD for all samples offline using both the fully trained checkpoint and an early checkpoint of the experimental model. Then, we develop a preference function to predict the model’s preference for data with specific PD characteristics, which allows us to pre-organize data based on the model’s preferences at different training stages. Because the data is fully arranged offline, the training of the experimental model proceeds without interruption.
The cost of calculating PD for the entire dataset is prohibitively high.
@@
Data with high PD are challenging for weak models but well-suited for strong models due to the increased capacity. Conversely, data with low PD are less sensitive to model capability differences and can be understood by both strong and weak models. From a FLOP perspective, the early stages of training an experimental model can be seen as involving a weak, smaller model, while later stages involve a strong, larger model. Thus, placing high-PD data in the later training stages can enhance the model’s ability to fit these challenging samples, whereas low-PD data can be trained earlier since they are less sensitive to model capability. This establishes a natural curriculum learning principle for the pretraining: begin with low-PD data and progress to high-PD data. Following the principle, we propose an S-shape function effectively that models preferences throughout training. Further, to maintain diversity in each batch, we use the concentration of high-PD data rather than PD itself as the output of the preference function.
By overcoming the challenges, PDPC can always perceive and use the data preferred by LLMs to pretrain and boost them. Notably, PDPC only arranges the given data without making other changes. In summary, our work has the following contributions: (1) We propose the PD to measure the difference in the fit of strong and weak models to samples, pointing out that high-PD data is challenging for weak models and is suitable to be arranged in the later pretraining stage. (2) We propose a novel PDPC framework that always perceives and uses the data preferred by LLMs to train and boost them, and ensures uninterrupted continuous training, which serves as the last data preprocessing step. (3) Experimental results show significant performance improvements over baselines. Notably, the 3B model trained on 1T tokens with PDPC demonstrates an average accuracy increase of 4.1% across all benchmarks and 8.1% across MMLU and CMMLU.
@@ 2 PD-based Preference Curriculum
We propose the PD-based Preference Curriculum learning framework, PDPC, which always perceives and uses the data preferred by LLMs to pretrain and boost them. PDPC is used as the last key data preprocessing step before pretraining. It only arranges the given data without other changes.
@@ 2.1 Problem Formulation
Given a pretraining dataset $\mathcal{D}$ with an initial random distribution, we aim to transform it into one that closely approximates the oracle distribution $\mathcal{O}_{\{B_{k}\}}$, which represents the ideal data distribution preferred by the model. To achieve this, we aim to find the optimal preference function $f^{*}$ that adjusts the joint distribution of all batches $\{B_{k}\}_{k=1}^{K}$ to closely approximate $\mathcal{O}_{\{B_{k}\}}$.
@@ 2.2 PD-based Data Partitioning
PD indicates the extent to which the strong model outperforms the weak model on a given sample. When the PD value is small, it indicates that the strong model $M_{s}$ and the weak model $M_{w}$ have a similar level of fit to the sample $x$. Conversely, a large PD value suggests that $M_{s}$ outperforms $M_{w}$ in fitting the sample $x$, implying that the sample is more challenging for $M_{w}$ to learn. In fact, there are very few samples whose perplexity in $M_{w}$ is smaller than in $M_{s}$, accounting for less than 0.01%, and we ignore these samples.
@@
An intuitive method for data reorganization is to sort the pretraining data by PD values from low to high. However, this method is suboptimal. Figure 6 and 7(b) show that high and low PD values differ significantly. Sorting by PD creates batches with overly homogeneous samples, lacking the diversity essential for pretraining LLMs Sachdeva et al. (2024). To solve this issue, we propose to divide data into distinct domains based on mutually exclusive PD ranges. During each pretraining step, data from each domain is mixed in varying proportions, enhancing both the effectiveness and diversity of the training process.
Formally, given a pretraining dataset $\mathcal{D}$ composed of multiple text samples, we calculate the PD for each sample $x$, denoted as $\text{PD}_{x}$. We then sort $\mathcal{D}$ in ascending order based on $\text{PD}_{x}$ and divide it evenly into $n$ sub-domains, $\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{n}$:
@@ 2.3 CL with PD-based Preference Function
We propose using a preference function $f(p)$ to model the model’s preferences for different PD partitions at any given training stage. The function maps the pretraining progress $p$ to the concentration of each domain, denoted as $\alpha_{i}$. The training completion rate $p$ is defined as: $p=\frac{k}{K}$, where $k$ is the current number of training steps and $K$ is the total number of training steps. The function $f(p)$ is defined as
$f$ $:[0,1]\rightarrow\mathbb{R}^{n}$ (5)
$f$ $(p)\rightarrow[\alpha_{1}(p),\alpha_{2}(p),\ldots,\alpha_{n}(p)]$
@@
which determines how the concentration of each domain changes as training progresses.
@@
Following the CL principle discussed in Section 1, we start with low-PD data and gradually move to high-PD data. To implement this, we divide the pretraining data into $n$ PD partitions, allowing us to explore the model’s preference for mixing ratios of data with different PD values at various pretraining stages. By understanding these preferences, we can tailor the curriculum learning approach to optimize model pretraining. We discuss different scenarios based on the number of domains $n$:
(1) When $n=1$, the training process involves randomly selecting individual samples from the entire dataset for each training step.
(2) When ${n=|\mathcal{D}|}$, each domain corresponds to a single sample, which is equivalent to performing a full sample-level sorting of all the data.
(3) When ${1<n<|\mathcal{D}|}$, each training step will include data from different domains. Finding the optimal weight function $f(r)$ is indeed a complex problem, often requiring significant time and resources to solve.
This paper focuses on the case where $n=2$ because it effectively captures the essential differences between low-PD and high-PD data while remaining computationally manageable. Essentially, $n=2$ divides the data into PD high and PD low parts. We introduce a PD concentration preference function to guide the offline organization of pretraining data. This function predicts the concentration of low PD data that the model prefers at different stages of training. However, directly optimizing the preference function $f$ is challenging. To solve this issue, we propose a function search method and discuss an annealing-based iterative optimization approach in Appendix B to approximate $f^{*}$. Specifically, to maintain a constant data volume and align with these preferences, it’s essential to choose the right function, which ensures that the proportion $b$ of low PD data matches the pretraining completion $p$, expressed as $b=f(p)$, and must meet certain criteria:
@@
Secondly, when $n=2$, we essentially divide the pretraining data into two equal parts based on PD. To ensure that the total amount of data with high PD and low PD remains equal during pretraining, the function can be symmetric about the point $(0.5,0.5)$, satisfying $f(0.5+\Delta)=1-f(0.5-\Delta)$, where $\Delta\in(0,0.5)$.
@@
We initially searched for three representative forms of preference functions as candidates: the linear function $f_{L}(p)$, the Z-shape function $f_{Z}(p)$, and the S-shape function $f_{S}(p)$, as shown in the left part of Figure 2. $f_{L}(p)$ represents a steady decline, while $f_{Z}(p)$ indicates a sudden, distinct change. To introduce a balance between the two, we also incorporate an S-shape function form.
@@
Essentially, our PDPC framework (Figure 2(b)) clusters preferred data within each batch, effectively performing Max-Pooling on this data. In contrast, traditional methods (Figure 2(a)) dilute the benefits of high PD data with less favorable low PD data, similar to average pooling.
@@
1:Input: pretraining data $D$, all training steps $K$
@@
6:Divide $D$ into $n$ sub-domains $D_{1},\ldots,D_{n}$ based on PD
7:Explore preference function $f(p)$
8:for $k=0$ to $K-1$ do
9:     $p\leftarrow\frac{k}{K}$
10:     $\bm{\alpha}(p)\leftarrow f(p)$
11:     Sample $B_{k}$ from $D_{i}$ according to $\alpha_{i}(p)$
12:     Update $\theta_{k}$ using $B_{k}$
13:end for
@@ 3 Experiments
3.1 Setup
Setup
We trained two experimental models: a 1.3B model using 100B randomly selected tokens from the SlimPajama dataset Soboleva et al. (2023), and a 3B model on a bilingual dataset containing 1T tokens. The bilingual dataset comprises 500B tokens each of Chinese and English data, sourced from domains such as booksGao et al. (2020), blogsBaumgartner et al. (2020), patentsSharma et al. (2019), Common CrawlPenedo et al. (2024), and Wikipedia, similar to the Matrix datasetZhang et al. (2024a). We train 100M and 700M reference models (RMs) on i.i.d. subsets from SlimPajama (50B) for the 1.3B setting, and 100M and 1.3B RMs on i.i.d. subsets for the 3B setting, respectively. All models were trained using the Llama architectureTouvron et al. (2023) within the Megatron framework Shoeybi et al. (2019), utilizing the Adam optimizer. We set batch size to 480 and the context window length to 8192. The initial learning rate is set to 2e-4, with a warm-up phase of over 375M tokens. We adopt a cosine learning rate schedule and set weight decay to 0.1.
@@ 3.1 Setup
We employ the lm-evaluation-harness Gao et al. (2021) to measure the models’ performance on the following benchmarks: ARC-E Clark et al. (2018), ARC-C Clark et al. (2018), SciQ Welbl et al. (2017), HellaSwag Zellers et al. (2019) and PIQA Bisk et al. (2020), which include tasks like knowledge question answering and commonsense reasoning. For the 3B model, we add benchmarks like MMLU Hendrycks et al. (2020), CMMLULi et al. (2024a), and CEVALHuang et al. (2023), which cover multi-domain knowledge and complex reasoning tasks, presenting challenges absent in 1.3B models. We employ in-context learning for evaluation following Wettig et al. (2024). Standard accuracy is used as the final metric for all tasks.
Baseline
We compared several basic curriculum learning approaches: (1) Random: Each batch is randomly selected from the entire dataset, corresponding to the case of $n=1$ in our framework. (2) PPL: PPL directly measures how well a model fits the data. We use two 700M models, each trained separately on the SlimPajama and Matrix subsets, to annotate their corresponding data. (3) QuRating Wettig et al. (2024): We select the Education Value in QuRating as the difficulty indicator for curriculum learning. It is worth noting that, unlike the setting in the QuRating paper, we use 100B data instead of the full 260B data of Slimpajama. (4) Sequential: We perform a complete sorting based on PD, PPL, and Qurating, either in ascending or descending order.

### v2

@@ Abstract
Large language models (LLMs) generally utilize a consistent data distribution throughout the pretraining process. However, as the model’s capability improves, it is intuitive that its data preferences dynamically change, indicating the need for pretraining with different data at various training stages. To achieve it, we propose the Perplexity Difference (PD) based Preference Curriculum learning (PDPC) framework, which always perceives and uses the data preferred by LLMs to train and boost them. First, we introduce the PD metric to quantify the difference in how challenging a sample is for weak versus strong models. Samples with high PD are more challenging for weak models to learn and are more suitable to be arranged in the later stage of pretraining. Second, we propose the preference function to approximate and predict the data preference of the LLM at any training step, so as to complete the arrangement of the dataset offline and ensure continuous training without interruption. Experimental results on 1.3B and 3B models demonstrate that PDPC significantly surpasses baselines. Notably, the 3B model trained on 1T tokens achieves an increased average accuracy of over 8.1% across MMLU and CMMLU.
@@
Xuemiao Zhang1,4,  Liangyu Xu4,  Feiyu Duan2,4thanks: Equal contribution.,  Yongwei Zhou4,  Sirui Wang3,4thanks: Corresponding author.,  Rongxiang Weng4,  Jingang Wang4,  Xunliang Cai4 1 Peking University 2 Beihang University 3 Tsinghua University 4 Meituan zhangxuemiao@pku.edu.cnduanfeiyu@buaa.edu.cnywzhouphd2018@gmail.com {xuliangyu02, wangsirui, wangjingang02, caixunliang}@meituan.com
@@ 1 Introduction
Large language models (LLMs) have shown impressive performance on various tasks after being pretrained on vast amounts of data Touvron et al. (2023); Dubey et al. (2024); Liu et al. (2024). As LLMs undergo extensive pretraining, their capabilities steadily improve, which influences their performance and data preferences Yu et al. (2024). Existing methods of uniformly sampling data throughout the pretraining process are suboptimal because they overlook the model’s evolving data preferences Wettig et al. (2024); Abbas et al. (2023); Sachdeva et al. (2024).
Recently, some research has shifted the focus to data influence on model capability during pretraining Evans et al. (2024); Yu et al. (2024); Koh and Liang (2017); Ko et al. (2024). A typical method named MATES considers the changing data influence on models but requires interrupting the training process to select more preferred data based on the model’s current state, which disrupts training continuity and stability Yu et al. (2024). A similar issue arises in JEST Evans et al. (2024).
In this paper, we introduce perplexity Ziegel et al. (1976) difference (PD) to quantify the difference in how challenging a sample is for LLMs’ final and early checkpoints and use PD to gain a deeper understanding of model characteristics. We further propose a novel PD-based Preference Curriculum learning framework, PDPC, as shown in Figure 1. It perceives LLMs’ data preference at any training step and uses the preferred data to continuously pretrain LLMs without interruption, thereby boosting their performance. To this end, PDPC needs to solve three major challenges:
@@
Ideally, training would be paused at any training step to calculate PD using the model’s current state, allowing for data sampling from the preference distribution like MATES Yu et al. (2024). However, to ensure continuity and stability in pretraining, we propose an offline processing method to approximate the ideal dynamic preference adjustment. First, we calculate PD for all samples offline using both the fully trained checkpoint and an early checkpoint of the experimental model. Then, we develop a preference function to predict the model’s preference for data with specific PD characteristics, enabling the pre-organization of data according to the model’s preferences at various training stages. Because the data is fully arranged offline, the training of the experimental model proceeds without interruption.
Calculating PD for the entire dataset is prohibitively expensive.
@@
Data with high PD are challenging for weak models but well-suited for strong models due to the increased capacity. Conversely, data with low PD are less sensitive to model capability differences and can be understood by both strong and weak models. From the perspective of FLOPs, the early stages of training an experimental model can be seen as involving a weak, smaller model, while later stages involve a strong, larger model. Thus, placing high-PD data in the later training stages can enhance the model’s ability to fit these challenging samples, whereas low-PD data can be trained earlier since they are less sensitive to model capability. This establishes a natural curriculum learning principle for the pretraining: begin with low-PD data and progress to high-PD data. Following the principle, we propose an S-shape function that effectively models preferences throughout training. Further, to maintain diversity in each batch, we use the concentration of low-PD data rather than PD itself as the output of the preference function.
By overcoming the challenges, PDPC can always perceive and use the data preferred by LLMs to pretrain and boost them. Notably, PDPC only arranges the given data without performing data selection. In summary, our work has the following contributions: (1) We propose PD to measure the difference in the fit of strong and weak models to samples, pointing out that high-PD data is challenging for weak models and is suitable to be arranged in the later pretraining stage. (2) We propose a novel PDPC framework that always perceives and uses the data preferred by LLMs to train and boost them, and ensures uninterrupted continuous training, which serves as the last data preprocessing step. (3) Experimental results show significant performance improvements over baselines. Notably, the 3B model trained on 1T tokens with PDPC demonstrates an average accuracy increase of 4.1% across all benchmarks and 8.1% across MMLU and CMMLU.
@@ 2 PD-based Preference Curriculum
In this section, we propose PDPC, which always perceives and uses the data preferred by LLMs to pretrain and boost them. Notably, PDPC only organizes the given data without performing selection, which can serve as the final data preprocessing step before pretraining.
@@ 2.1 Problem Formulation
Given a pretraining dataset $\mathcal{D}$ following a uniform distribution, we aim to arrange it into a form that better aligns with the oracle distribution $\mathcal{O}_{\{B_{k}\}}$, which represents the ideal data distribution preferred by the model. To achieve this, we aim to find the optimal preference function $f^{*}$ that adjusts the joint distribution of all batches $\{B_{k}\}_{k=1}^{K}$ to closely approximate $\mathcal{O}_{\{B_{k}\}}$.
@@ 2.2 PD-based Data Partitioning
PD indicates the extent to which the strong model outperforms the weak model on a given sample. Low PD indicates that the strong model $M_{s}$ and the weak model $M_{w}$ have a similar level of fit to the sample $x$. Conversely, a high PD value suggests that $M_{s}$ outperforms $M_{w}$ in fitting the sample $x$, implying that the sample is more challenging for $M_{w}$ to learn. There are very few samples whose perplexity in $M_{w}$ is smaller than in $M_{s}$, accounting for less than 0.01%, and we ignore them.
@@
An intuitive method for data arrangement is to sort the pretraining data by PD from low to high. However, this method is suboptimal. As shown in Figure 6 and 7 of Section 3.5, data with extremely high and extremely low PD values differ significantly. Sorting by PD creates batches with overly homogeneous samples, lacking the diversity essential for pretraining LLMs Sachdeva et al. (2024).
To solve this issue, we propose to partition data into distinct parts based on mutually exclusive PD ranges. During each pretraining step, data from each part is mixed in varying proportions. Formally, given a pretraining dataset $\mathcal{D}$, we calculate the PD for each data point $x$, denoted as $\text{PD}_{x}$. We then sort $\mathcal{D}$ in ascending order based on $\text{PD}_{x}$ and partition it evenly into $n$ parts, $\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{n}$:
@@ 2.3 CL with PD-based Preference Function
We introduce a preference function $f(p)$ to capture the model’s preferences for different PD partitions at any training step. The function maps the pretraining progress $p$ to a proportion vector $\bm{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{n}]$, where $\alpha_{i}$ denotes the fraction of data from the $i$-th domain in the current batch. The training completion rate $p$ is defined as: $p=\frac{k}{K}$, where $k$ is the current training step and $K$ is the total number of training steps. The function $f(p)$ is defined as
$f$ $(p)\rightarrow[\alpha_{1}(p),\alpha_{2}(p),\ldots,\alpha_{n}(p)]$ (5)
@@
which determines how the proportion of each part changes as training progresses.
@@
Following the CL principle discussed in Section 1, PDPC starts the training process using low-PD data and gradually moves to high-PD data. To implement this, we sort the pretraining data and partition them into $n$ parts, allowing us to explore the model’s preference for mixing ratios of data with different PD values at various pretraining steps. We discuss different scenarios:
(1) When $n=1$, the training process involves randomly selecting individual samples from the dataset for each training step.
(2) When ${n=|\mathcal{D}|}$, each part corresponds to a single sample, which is equivalent to performing a full sample-level sorting of all the data.
(3) When ${1<n<|\mathcal{D}|}$, each training step will include data from different parts. Finding the optimal function $f^{\star}$ is a complex problem that involves substantial costs.
We focus on the case where $n=2$ because it effectively captures the essential differences between low-PD and high-PD data while remaining computationally manageable. Essentially, $n=2$ partitions the data into high-PD and low-PD parts.
To guide the offline organization of pretraining data, We introduce a PD-based preference function $f(\frac{k}{K})$, which predicts the proportion of low PD data that the model prefers at different training steps $k$. However, directly optimizing the preference function $f$ is challenging. To solve this issue, we propose a function search method [^1] We also discuss an annealing-based iterative optimization approach in Appendix B to approximate $f^{*}$. to approximate $f^{*}$. To ensure equal volumes of high and low PD data and align with the model’s data preferences, it is crucial to choose the right function. This function ensures that the proportion $b$ of low PD data matches the pretraining completion $p=\frac{k}{K}$, expressed as $b=f(p)$, and must meet specific criteria:
@@
Secondly, to ensure that the total amount of data with high PD and low PD stays equal, the function can be symmetric about the point $(0.5,0.5)$, satisfying $f(0.5+\Delta)=1-f(0.5-\Delta)$, where $\Delta\in(0,0.5)$.
@@
We initially search for three representative functions as candidates: the linear function $f_{L}(p)$, the Z-shape function $f_{Z}(p)$, and the S-shape function $f_{S}(p)$, as shown in the left part of Figure 2. $f_{L}(p)$ represents a steady decline, while $f_{Z}(p)$ indicates a sudden, distinct change. To introduce a balance between the two, we also incorporate an S-shape function form.
@@
Essentially, assuming that at step $i$ the model prefers high PD data, traditional methods (Figure 2(a)) dilute the benefits of high PD data with less favorable low PD data, similar to average pooling. In contrast, PDPC (Figure 2(b)) clusters preferred data within each batch, resembling the effect of Max-Pooling on this data.
@@
1:Input: dataset $\mathcal{D}$, total iterations $K$, batch size $N$
@@
6:Partition $\mathcal{D}$ into $2$ sub-domains $A_{PD}^{low}$ and $A_{PD}^{high}$.
7:Explore and determine the form of preference function:
8:  $f(p)=\frac{1}{1+exp{(a(p-0.5))}}$
9:for $k=0$ to $K-1$ do
10:     Calculate pretraining progress $p=\frac{k}{K}$
11:     Get the proportions vector:
12:      $[\alpha_{1},\alpha_{2}]\leftarrow[f(p),1-f(p)]$
13:     Sample from the two domains to form $B_{k}$:
14:     $B_{k}=\{x\mid x\sim A_{PD}^{low}\}_{\alpha_{1}\cdot N}\cup\{x\mid x\sim A_{PD}^{high}\}_{\alpha_{2}\cdot N}$
15:     Train the model on $B_{k}$ and update $\theta_{k}$
16:end for
@@ 3 Experiments
3.1 Experimental Setup
General Setting
We train two experimental models: a 1.3B model using 100B randomly selected tokens from the SlimPajama dataset Soboleva et al. (2023), and a 3B model on a bilingual dataset containing 1T tokens, which comprises 500B tokens each of Chinese and English data, sourced from domains such as booksGao et al. (2020), blogsBaumgartner et al. (2020), patentsSharma et al. (2019), Common CrawlPenedo et al. (2024), and Wikipedia, similar to the Matrix datasetZhang et al. (2024a). We train 100M and 700M reference models (RMs) on an i.i.d. subset with 50B tokens from SlimPajama for the 1.3B setting, and 100M and 1.3B RMs on an i.i.d. subset with 500B tokens for the 3B setting, respectively. All models were trained using the Llama architectureTouvron et al. (2023) within the Megatron framework Shoeybi et al. (2019), utilizing the Adam optimizer. We set the batch size to 640 and the context window length to 8192. The initial learning rate is set to 2e-4, with a warm-up phase of over 375M tokens. We adopt a cosine learning rate schedule and set weight decay to 0.1. A full pretraining run of the 3B model on 1 trillion tokens, utilizing 512 Ascend 910B NPUs, requires approximately 180 hours.
@@ 3.1 Experimental Setup
We employ the lm-evaluation-harness Gao et al. (2021) to measure the models’ performance on the following benchmarks: ARC-E Clark et al. (2018), ARC-C Clark et al. (2018), SciQ Welbl et al. (2017), HellaSwag Zellers et al. (2019) and PIQA Bisk et al. (2020), which include tasks like knowledge question answering and commonsense reasoning. For the 3B model, we add benchmarks like MMLU Hendrycks et al. (2020), CMMLULi et al. (2024), and CEVALHuang et al. (2023), which cover multi-domain knowledge and complex reasoning tasks, presenting challenges absent in 1.3B models. We employ in-context learning for evaluation following QuRating Wettig et al. (2024). Standard accuracy is used as the final metric for all tasks.
Baselines
We compare PDPC with Random and several basic curriculum learning approaches: (1) Random: Each batch is randomly selected from the entire dataset, corresponding to the case of $n=1$ in our framework. (2) PPL: PPL directly measures how well a model fits the data. We use two 700M models, each trained separately on the SlimPajama and Matrix subsets, to annotate their corresponding data. (3) QuRating Wettig et al. (2024): We select the Education Value in QuRating as the difficulty indicator for curriculum learning. (4) Sequential: We fully sort the data based on PD, PPL, and Qurating, arranging them in either ascending or descending order.